ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 199, 2014

Alon, N., J. Balogh, R. Morris and W. Samotij Counting sumfree sets in abelian groups

309

Ballester-Bolinches, A., L. M. Ezquerro and A. N. Skiba On subgroups of hypercentral type of finite groups

259

Balogh, J., R. Morris and W. Samotij Random sum-free subsets of abelian groups

651

Balogh, J. See N. Alon Becher, K. J., D. B. Leep and C. Schubert Semiorderings and

stability index under field extensions

547

Bernstein, J. and B. Krötz Smooth Fréchet globalizations of Harish-Chandra modules

45

Bertolini, M. and H. Darmon Kato’s Euler system and rational points on elliptic curves I: A p-adic Beilinson formula

163

Burness, T. C., R. M. Guralnick and J. Saxl Base sizes for S-actions of finite classical groups

711

Carando, D. and P. Sevilla-Peris Extendibility of bilinear forms on banach sequence spaces

941

Caselli, F. and E. Marberg Isomorphisms, automorphisms, and generalized involution models of projective reflection groups

433

Černe M. and M. Flores On Beurling’s boundary differential relation

831

Cominetti, R., J. A. Soto and J. Vaisman On the rate of convergence of Krasnosel’skiǐ–Mann iterations and their connection with sums of Bernoullis

757

Dankwart, K. Rigidity of flat surfaces under the boundary measure  623

Darmon, H. See M. Bertolini Deng, S. and M. Xu Clifford–Wolf translations of homogeneous Ran-

ders spheres

507

Dicks, W. and C. Martínez-Pérez Isomorphisms of Brin–Higman– Thompson groups

189

Ezquerro, L. M. See A. Ballester-Bolinches Felikson, A. and P. Tumarkin Essential hyperbolic Coxeter poly-

topes

113

Flores, M. See M. Černe

Goldbring, I. and H. Towsner An approximate logic for measures  867

Guralnick, R. M. See T. C. Burness

Hujter, M. and Z. Lángi On the multiple Borsuk numbers of sets  219
Iovanov, M. C. On extensions of rational modules  585

Jaksztas, L. On the left side derivative of the Hausdorff dimension of the Julia sets for z^{2} + c at c = −3/4

1

Jeon, W., I. Kim, K. Ohshika and C. Lecuire Primitive stable representations of free Kleinian groups

841

Keller, T. M. and Y. Yang Orbits of finite solvable groups on characters

933

Kim, I. See W. Jeon Kosloff, Z. On the K property for Maharam extensions of Bernoulli

shifts and a question of Krengel

485

Krötz, B. See J. Bernstein Kucherenko, T. and C. Wolf Geometry and entropy of generalized

rotation sets

791

Kumar, M. Killing wild ramification  421
Kwaśniewski, B. K. Topological freeness for Hilbert bimodules  641

Lángi, Z. See M. Hujter Lecuire, C. See W. Jeon Leep, D. B. See K. J. Becher Lindenstrauss, E. and M. Tsukamoto Mean dimension and an

embedding problem: an example

573

Louder, L. Simple loop conjecture for limit groups  527

Luiro, H. On the Hamilton–Jacobi equation and infimal convolution in the framework of Sobolev-functions

267

Marberg, E. See F. Caselli Martínez-Pérez, C. See W. Dicks Matter, M. and T. Nagnibeda The abelian sandpile model on ran-

domly rooted graphs and self-similar groups

363

Mishchenko, S. and A. Valenti An almost nilpotent variety of exponent 2

241

Morris, R. See N. Alon Morris, R. See J. Balogh Nagnibeda, T. See M. Matter Oh, H. and N. A. Shah Limits of translates of divergent geodesics

and integral points on one-sheeted hyperboloids

915

Ohshika, K. See W. Jeon Peres, Y., J. Schmeling, S. Seuret and B. Solomyak Dimen-

sions of some fractals defined via the semigroup generated by 2 and 687 Ramana, D. S. and O. Ramaré Additive energy of dense sets of

primes and monochromatic sums

955

Ramaré, O. See D. S. Ramana

Rinot, A. The Ostaszewski square and homogeneous Souslin trees  975

Samotij, W. See N. Alon Samotij, W. See J. Balogh Saxl, J. See T. C. Burness Schmeling, J. See Y. Peres

Schoen, T. and I. D. Shkredov Roth’s theorem in many variables  287

Schubert, C. See K. J. Becher Seuret, S. See Y. Peres Sevilla-Peris, P. See D. Carando Shah, N. A. See H. Oh Shelah, S. and L. Strüngmann On the p-rank of Ext(A,B) for

countable abelian groups A and B

567

Shen, Z. and G. Yang On a class of weakly Einstein Finsler metrics  773

Shkredov, I. D. See T. Schoen Skiba, A. N. See A. Ballester-Bolinches Solomyak, B. See Y. Peres Soto, J. A. See R. Cominetti Strüngmann, L. See S. Shelah Towsner, H. See I. Goldbring Tsukamoto, M. See E. Lindenstrauss Tumarkin, P. See A. Felikson Vaisman, J. See R. Cominetti Valenti, A. See S. Mishchenko Wolf, C. See T. Kucherenko Xu, M. See S. Deng Yang, G. See Z. Shen

Yang, Y. Large orbits of subgroups of solvable linear groups  345

Yang, Y. See T. M. Keller